Carnes Cosmetics Co.'s stock price is $74.81, and it recently paid a $1.75 dividend. This dividend is expected to grow by 27% for the next 3 years, then grow forever at a constant rate, g; and rs = 13%. At what constant rate is the stock expected to grow after Year 3? Round your answer to two decimal places. Do not round your intermediate calculations.
Answer #1

Price = P0= 74.81

D0 = 1.75

Growth = 27% for 3 years

rs=13%

Hence Dividends are

D1= 1.75*127% = 2.2225

D2= 2.2225* 127% = 2.822575

D3= 2.822575* 127%= 3.58467

D4= 3.58467* 127% = 4.552531

Stock price is the discounted dividends + Horizon value

Horizon value = P3 = D4/(rs-g)

Hence

74.81 = 2.2225/ 1.13^1 + 2.822575/1.13^2 + 3.58467/1.13^3 + 4.552531/ (0.13-g)

74.81=6.661661+ 4.552531/ (0.13-g)

4.552531/ (0.13-g)= 68.14834

0.13-g= 0.066803

g=0.063197

Growth = 6.32%